An oxygen gas container has a volume of 20.0 L. How many grams of oxygen are in the container if the gas has a pressure of 887 mmHg at 24 ∘C?

Mg metal reacts with HCl to produce hydrogen gas.
Mg(s)+2HCl(aq)→MgCl2(aq)+H2(g)

What volume of hydrogen at 0 ∘C and 1.00 atm (STP) is released when 9.40 g of Mg reacts?

How many grams of magnesium are needed to prepare 6.90 L of H2 at 775 mmHg and 20 ∘C?
Answer #1

1)

Given:

P = 887 mm Hg

= (887/760) atm

= 1.1671 atm

V = 20.0 L

T = 24.0 oC

= (24.0+273) K

= 297 K

find number of moles using:

P * V = n*R*T

1.1671 atm * 20 L = n * 0.08206 atm.L/mol.K * 297 K

n = 0.9577 mol

Molar mass of O2 = 32 g/mol

use:

mass of O2,

m = number of mol * molar mass

= 0.9577 mol * 32 g/mol

= 30.65 g

Answer: 30.6 g

Only 1 question at a time please